(5 x 2 = 10 pts) Consider the following functions. For each of them, determine how many times is ‘hi printed in terms of the1 2 def fun (n) { if (n > 1) { for i=1 to n { print( hihi ) 3 4 5 ها 6 fun(n/3) fun (n/3) 7 00 } 9}
Answer #1

a> T(n) = 27 (113) + constant 1 T(n) ty) Tin/3) +(1937) (1932) TM/3 -T(%3k) 11 Time complexity = 2 KAT -1 - 35 k= logan na =Tn)= n 2T | 3) + cost n n 20/3 2² n/3- /3 2/37 /3.5 /32 /k ak n/zk K=1093n #levels - Total #callo = kn+ 2 TC Max) logan n